OpenStudy (anonymous):

Please help with this. @Sionainn @jdoe0001 @jdoe0001

OpenStudy (jdoe0001):

well.. is the same another i = prt and you'd need to solve for "r" notice the graph 

OpenStudy (jdoe0001):

at 0 years. the amount starts off with 326 so. that's the original deposit amount and thus that's the Principal the other dots tell how much is in the account at those years so if you subtract the "A"mount or balance from the "P"rincipal... that'd give the "interest accrued" and then plug them in the solved equation for "r"

OpenStudy (jdoe0001):

we.... subtract the Principal from the Balance.. anyhow.. .that'd be the "interest accrued"

OpenStudy (anonymous):

357.25-326= interest? @jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

x=0.019 x=1.9% Final Answer? @jdoe0001

OpenStudy (jdoe0001):

yeap 0.01917177914110429448 * 100 = 1.9%

OpenStudy (anonymous):

I thought you just move the decimal over twice...., can't you ?

OpenStudy (jdoe0001):

yeap.. you do \(\bf 0.01917177914110429448\times 100\to 00\underline{1.9}17177914110429448 \\ \quad \\ \underline{1.9}17177914110429448 \approx \underline{1.9}\%\)

OpenStudy (jdoe0001):

or 1.92% if you wish.. or 1.9172%